# HyperPg - Prototypical Gaussians on the Hypersphere for Interpretable Deep Learning

### v2

@@ Abstract
Prototype Learning methods provide an interpretable alternative to black-box deep learning models. Approaches such as ProtoPNet learn, which part of a test image “look like” known prototypical parts from training images, combining predictive power with the inherent interpretability of case-based reasoning. However, existing approaches have two main drawbacks: A) They rely solely on deterministic similarity scores without statistical confidence. B) The prototypes are learned in a black-box manner without human input. This work introduces HyperPg, a new prototype representation leveraging Gaussian distributions on a hypersphere in latent space, with learnable mean and variance. HyperPg prototypes adapt to the spread of clusters in the latent space and output likelihood scores. The new architecture, HyperPgNet, leverages HyperPg to learn prototypes aligned with human concepts from pixel-level annotations. Consequently, each prototype represents a specific concept such as color, image texture, or part of the image subject. A concept extraction pipeline built on foundation models provides pixel-level annotations, significantly reducing human labeling effort. Experiments on CUB-200-2011 and Stanford Cars datasets demonstrate that HyperPgNet outperforms other prototype learning architectures while using fewer parameters and training steps. Additionally, the concept-aligned HyperPg prototypes are learned transparently, enhancing model interpretability.
@@ 1 Introduction
Deep Learning has achieved high accuracy in many computer vision tasks. However, the decision-making processes of these models lack transparency and interpretability, making deployment in safety-critical areas challenging. Explainable Artificial Intelligence (XAI) seeks to develop interpretability methods to open the black-box reasoning processes of these models and increase trust in their decisions.
@@
XAI methods can be broadly divided into two categories: First, Post-Hoc Methods like LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017) or GradCAM Selvaraju et al. (2017) offer explanations for predictions without requiring retraining. While applicable in many scenarios, post-hoc methods may not actually align with the models’ decision making processes, potentially leading to interpretations that are not entirely faithful (Rudin, 2019). Second, inherently interpretable methods provide built-in, case-based reasoning processes. For instance, small decision trees are inherently interpretable because their reasoning can be easily understood as a series of if-else statements (Molnar, 2020). However, they are constrained in their representational power.
@@
Deep Prototype Learning Architectures such as ProtoPNet (Chen et al., 2019) and its derivatives (e.g., Rymarczyk et al., 2020; Donnelly et al., 2021; Sacha et al., 2023) integrate inherent interpretability into deep learning models through a prototype layer. Each neuron in this layer represents a prototype, storing a latent feature vector. The model’s predictions are based on the distances between sample features and prototype parameters, for example by computing the $L_{2}$-distance. However, these deterministic similarity scores do not include statistical information like confidence. To include such contextual information, prototypes could be modeled as probability distributions, like Gaussian distributions. Yet, recent prototype architectures favor the cosine hypersphere as feature space for its classification advantages (Mettes et al., 2019), where defining a Gaussian distribution analogue is challenging (Hillen et al., 2017).
The prototypes in ProtoPNet and its successors are learned in an opaque manner via backpropagation. After training, the model “pushes” the learned prototypes on the embedding of the most similar known training sample (Chen et al., 2019). However, these prototypes are learned without human control and do not incorporate domain knowledge. A new training regime is needed to aligned prototypes with human-defined concepts. The main contributions of this paper are as follows:
@@
  • HyperPg: A new prototype representation with learned parameters anchor $\alpha$, mean $\mu$ and standard deviation $\sigma$. This representation models a Gaussian distribution over cosine similarities, thereby projecting a Gaussian distribution on the surface of a hypersphere. HyperPg’s similarity score is based on the Gaussian’s probability density function and adapts its size through a learned standard deviation.
  • HyperPgNet: A new prototype learning architecture built on HyperPg prototypes. HyperPgNet learns prototypes aligned with human-defined concepts using pixel level annotations. These concepts describe features such as color and patterns of birds or car parts.
  • Concept Extraction Pipeline: Based on Grounding DINO (Liu et al., 2023) and SAM2 (Ravi et al., 2024), this pipeline provides pixel-level concept annotations at scale.
  • Classification Experiments: HyperPgNet is compared with other prototype learning architectures like ProtoPNet on the CUB-200-2011 (Wah et al., 2011) and Stanford Cars (Krause et al., 2013) datasets. HyperPgNet outperforms the other models with fewer learned prototypes, while infusing each prototype with more meaning.
@@ 2 Related Work
Prototype Learning. In image classification, prototype learning approaches using autoencoders provide high interpretability by reconstructing learned prototypes from latent space back to the image space (Li et al., 2018). However, these approaches are limited in their performance because each prototype must represent the entire image. ProtoPNet (Chen et al., 2019) introduced the idea of prototypical parts. In this setting, each prototype is a latent patch of the input image, commonly a $1\times 1$ latent patch. The prototypes are each associated with a single class and learned via backpropagation without additional information.
HyperPgNet builds on the idea of prototypical parts, but learns the prototypes in a transparent, concept-aligned manner. The prototypes are not class exclusive, but class shared. Each prototype corresponds to a human defined concept. To provide the required concept annotations, we propose a labeling pipeline based on foundation models such as DINOv2 (Oquab et al., 2024) and SAM2 (Ravi et al., 2024). HyperPgNet learns the concept-aligned prototypes by only adapting the loss functions and providing additional annotations, without requiring changes to the feature encoder like PIPNet (Nauta & Seifert, 2023) or Lucid-PPN (Pach et al., 2024), which uses a hybrid input head to disentangle prototype color and shape.
Multiple successors build on the idea of ProtoPNet. ProtoPShare (Rymarczyk et al., 2020) and ProtoPool (Rymarczyk et al., 2022) use additional optimizations to learn class-shared prototypical parts. Deformable ProtoPNet (Donnelly et al., 2021) learns a mixture of prototypical parts with dynamic spatial arrangement. Other work replace the linear output layer of ProtoPNet with other interpretable models: ProtoKNN (Ukai et al., 2023) uses a k-nearest neighbor classifier and ProtoTree (Nauta et al., 2021) employs a decision tree.
All these models change how the prototype activations are further processed. However, they all rely on the point-based prototype formulation introduced by ProtoPNet based on the $L_{2}$ or cosine similarity. HyperPgNet proposes a novel prototype formulation based on Gaussian distributions on the hypersphere, without changing the downstream processing. This probabilistic view with learned mean and variance allows HyperPgNet to learn prototypes with different degrees of specialization. Furthermore, the statistical confidence derived from the learned variance could be integrate in further downstream processing.
@@
The view of prototype learning as clustering in latent space was introduced for image segmentation (Zhou et al., 2022). The image patches are points in latent space with the prototypes as cluster centers. ProtoGMM (Moradinasab et al., 2024) and MGProto (Wang et al., 2023) build on this idea to model the prototypes as Gaussian mixture models for image segmentation.
HyperPg adapts the idea of Gaussian prototypes used in image segmentation to hyperspherical latent spaces based on the cosine similarity. The cosine similarity has been shown to perform well in classification tasks (Mettes et al., 2019) and also been applied in recent prototype learning works (Zhou et al., 2022; Ukai et al., 2023). This makes HyperPgNet the first model to employ probabilistic prototypes with learnable mean and variance in a hyperspherical space.
Concept-Based Methods. Testing with Concept Activation Vectors (TCAV) (Kim et al., 2018) is a post-hoc method that aligns model activations with human-understandable concepts. In image classification, TCAV measures model activations on images representing a specific concept like “striped” and compares them to activations on a random image set. Concept Activation Regions (CAR) (Crabbé & van der Schaar, 2022) relax TCAV’s assumption of linear separability in latent space by using support vector machines and the kernel trick. Another post-hoc approach, CRAFT (Fel et al., 2023), generates concept attribution maps by identifying relevant concepts from the training set via Non-Negative Matrix Factorization (NMF) and backpropagating with GradCAM (Selvaraju et al., 2017).
Instead of analyzing the learned concepts post-hoc, some approaches integrate concept learning into the classifier. MCPNet (Wang et al., 2024) extracts concepts from multiple layers by splitting feature maps along the channel dimension, learning concepts during training and at multiple scales. Concept Bottleneck Models (CBMs) (Koh et al., 2020) augment black-box models to predict pre-defined concept labels in an intermediate step. In this manner, CBMs behave like two sequential black box models and their internal reasoning remains opaque.
HyperPgNet combines the concept-aligned learning approach of CBM with interpretable prototype learning. HyperPgNet learns multiple HyperPg prototypes per concept, instead of a fully connected concept prediction layer. A novel concept extraction pipeline, built on foundation models, provides pixel level annotations at scale. By using those annotations to learn concept prototypes, HyperPgNet avoids the ambiguity of aligning post-hoc explanations with the model’s internal reasoning.
3 Prototypical Gaussians on the Hypersphere
@@
Figure 1: Illustration on how the different prototype formulations compute the similarity between a prototype ${\bm{p}}$ and latent vector ${\bm{z}}$. $L_{2}$ prototypes compute the Euclidean distance between two points in latent space. Hyperspherical prototypes use the cosine similarity of normalized vectors, which corresponds to the angle between two points on a hypersphere. Gaussian prototypes model a Gaussian distribution in Euclidean space and compute the probability density function (PDF). HyperPg prototypes learn a Gaussian distribution of cosine similarities, thereby projecting a Gaussian distribution onto the surface of a hypersphere.
@@
Prototype learning for deep neural networks involves finding structures in latent space representations. This section provides an overview of existing methods, which are also illustrated in Fig. 1. Prior work uses point-based prototypes, computing similarity scores relative to a single point in latent space. Aligning prototypes with human-labeled concepts requires a more powerful prototype representation, leading to the introduction of HyperPg - Prototypical Gaussians on the Hypersphere. HyperPg prototypes are able to adapt their shape in the latent space by modeling Gaussian distributions with mean and standard deviation, adapting to the variance in encoded concept features.
@@ 3.1 Point Based Prototypes
Autoencoder approaches use $\rho=\zeta$, meaning the prototype represents the entire image and can be reconstructed from latent space (Li et al., 2018). Part-based approaches like ProtoPNet and segmentation models like ProtoSeg use $\rho=1$ (Chen et al., 2019; Zhou et al., 2022), meaning each prototype represents some part of the image. Notable exceptions include Deformable ProtoPNet (Donnelly et al., 2021), where each prototype has spatial size $\rho=3\times 3$ and MCPNet (Wang et al., 2024), where prototypes are obtained by dividing the channels of the latent space into chunks.
The prediction is computed by comparing each prototype ${\bm{p}}$ to the latent feature map ${\bm{z}}$. For simplicity’s sake lets assume the spatial dimensions $\rho=\zeta=1$. The following equations can be adapted for higher spatial dimensions by summing over $\sum_{\rho_{w}}\sum_{\rho_{h}}$ for each chunk of the latent map.
ProtoPNet’s prototypes leverage the $L_{2}$ similarity. The $L_{2}$ similarity measure is defined as
@@
and is based on the inverted $L_{2}$ distance between a latent vector ${\bm{z}}$ and a prototype vector ${\bm{p}}$. This similarity is a point-based measure, as only the two vectors are compared, without any additional context like the expected variance of the cluster represented by the prototype.
Hyperspherical prototypes using the cosine similarity have been shown to perform well in classification tasks (Mettes et al., 2019) and have been widely used since (e.g., Zhou et al., 2022; Ukai et al., 2023). The cosine similarity is defined as
@@
which is based on the angle between two normalized vectors of unit length. By normalizing $D$ dimensional vectors to unit length, they are projected onto the surface of a $D$ dimensional hypershere. The cosine similarity is defined on the interval $[-1,1]$ and measures: $1$ for two vectors pointing in the same direction, 00 for orthogonal vectors, and $-1$ for vectors pointing in opposite directions. Like the $L_{2}$ similarity, the cosine similarity is a point-based measure comparing only two vectors.
Both the $L_{2}$ and cosine similarity can be used for classification. The similarity scores are processed by a fully connected layer (e.g., Chen et al., 2019; Donnelly et al., 2021), or a winner-takes-all approach assigns the class of the most similar prototype (e.g., Sacha et al., 2023). Prototypes can be learned by optimizing a task-specific loss, such as cross-entropy, via backpropagation. Alternatively, Zhou et al. (2022) propose “non-learnable” prototypes, whose parameters are obtained via clustering in the latent space rather than backpropagation.
3.2 Gaussian Prototypes
Gaussian prototypes model prototypes as a Gaussian distribution with mean and covariance. They adapt to the spread of the associated latent cluster by adjusting their covariance matrix. Thus, a Gaussian prototype with a wide covariance can still have a relatively high response even for larger distances from the mean vector.
Let the formal definition of a Gaussian prototype be ${\bm{p}}^{G}=(\bm{{\mu}},\bm{\mathrm{\Sigma}})$. The parameters of ${\bm{p}}^{G}_{c,j}$ now track both the mean and covariance of latent vector distribution ${\bm{Z}}_{c,j}$. Each Gaussian prototype ${\bm{p}}^{G}$ thus defines a multivariate Gaussian Distribution ${\mathcal{N}}(\bm{{\mu}},\bm{\mathrm{\Sigma}})$. Gaussian prototypes can be trained using EM for clustering in the latent space (Zhou et al., 2022; Wang et al., 2023; Moradinasab et al., 2024) or by directly optimizing the parameters via Backpropagation. The similarity measure of Gaussian prototypes is defined as the probability density function (PDF) for $D$-dimensional multivariate Gaussians, namely

### v3

# An Overview of Prototype Formulations for Interpretable Deep Learning

@@ Abstract
Prototypical part networks offer interpretable alternatives to black-box deep learning models. However, many of these networks rely on Euclidean prototypes, which may limit their flexibility. This work provides a comprehensive overview of various prototype formulations. Experiments conducted on the CUB-200-2011, Stanford Cars, and Oxford Flowers datasets demonstrate the effectiveness and versatility of these different formulations.
XAI, Prototype Learning, Image Classification, Deep Learning
@@ 1 Introduction
Deep Learning has achieved high accuracy in many computer vision tasks. However, the decision-making processes of these models lack transparency and interpretability, making deployment in safety-critical areas challenging. Explainable Artificial Intelligence (XAI) seeks to develop interpretability methods to open the black-box reasoning processes of such deep models and increase trust in their decisions.
@@
XAI methods can be broadly divided into two categories: First, Post-Hoc Methods like LIME (Ribeiro et al., 2016), SHAP (Lundberg & Lee, 2017) or GradCAM (Selvaraju et al., 2017) offer explanations for predictions without requiring retraining. While applicable in many scenarios, post-hoc methods may not actually align with the models’ decision-making processes, potentially leading to interpretations that are not entirely faithful (Rudin, 2019). Second, inherently interpretable methods provide built-in, case-based reasoning processes. For instance, small decision trees are inherently interpretable because their reasoning can be easily understood as a series of if-else statements (Molnar, 2020). However, they are constrained in their representational power.
Figure 1: Different Prototype Formulations. HyperPG is a novel formulation for probabilistic prototypes on a Hypersphere.
@@
Deep Prototype Learning Architectures such as ProtoPNet (Chen et al., 2019) and its derivatives (e.g., Rymarczyk et al., 2020; Donnelly et al., 2021; Sacha et al., 2023) integrate inherent interpretability into deep learning models through a prototype layer. Each neuron in this layer represents a prototype, storing a latent feature vector. The model’s predictions are based on the distances between sample features and prototype parameters, for example, by computing the $L_{2}$-distance. This makes prototype learning essentially a clustering task in latent space (Zhou et al., 2022). Different prototype formulations, such as prototypes using the cosine similarity, induce a hyperspherical structure to the latent space with performance advantages to classification (Mettes et al., 2019). Probabilistic formulations such as Gaussian prototypes enable further downstream tasks relying on probability values such as Bayesian approaches.
With the increased popularity of prototypical part models, there has been a corresponding rise in the number of architectural choices. However, to the best of our knowledge, there is no comprehensive overview of the design choices and prototype formulations for learning interpretable prototypes and their effect on predictive performance.
Table 1: Overview of 14 related models and their prototype learning configurations. 8/14 models use Euclidean prototypes.
Model Similarity Shape Assignment Clf. Head
Prototype Decoder Euclidean Entire Image Class Exclusive FCL Li et al. (2018)
ProtoPNet Euclidean Patch Class Exclusive FCL Chen et al. (2019)
Def. ProtoPNet Cosine Spatial Arrangement Class Exclusive FCL Donnelly et al. (2021)
ProtoPShare Euclidean Patch Merged after Training FCL Rymarczyk et al. (2020)
ProtoPool Euclidean Patch Pooling Shared FCL Rymarczyk et al. (2022)
TesNet Grassman Patch Class Exclusive FCL Wang et al. (2021)
ProtoTree Euclidean Patch Shared Decision Tree Nauta et al. (2021)
ProtoKNN Cosine Patch Class Exclusive KNN Clf Ukai et al. (2023)
PIPNet Cosine Patch Shared FCL Nauta et al. (2023)
LucidPPN Euclidean Separate Color & Texture Class Exclusive Branch Aggregation Pach et al. (2024)
ProtoSeg Cosine Patch Class Exclusive FCL Zhou et al. (2022)
ProtoGMM Gaussian Patch Class Exclusive FCL Moradinasab et al. (2024)
MGProto Gaussian Patch Class Exclusive Bayesian Likelihood Wang et al. (2024b)
ProtoPFormer Euclidean Transformer Token Class Exclusive Branch Aggregration Xue et al. (2024)
This paper explores different prototype formulations. Its contributions are as follows:
Prototype Formulation Overview: A comprehensive overview of different, learnable prototype definitions, including both point based and probabilistic prototypes as shown in Figure 1.
@@
HyperPG: A new probabilistic prototype representation with learned parameters anchor $\alpha$, mean $\mu$ and variance $\sigma^{2}$. This representation models a Gaussian distribution over cosine similarities, thereby projecting a Gaussian distribution on the surface of a hypersphere. HyperPG’s similarity score is based on the Gaussian’s probability density function and adapts its size through a learned standard deviation.
Benchmarking Protoypes: Extensive image classification experiments based on prototypical part network architectures like ProtoPNet on the CUB-200-2011 (Wah et al., 2011), Stanford Cars (Krause et al., 2013) and Oxford Flowers (Nilsback & Zisserman, 2008) datasets.
@@ 2 Related Work
Prototype Learning. In image classification, prototype learning approaches using autoencoders provide high interpretability by reconstructing learned prototypes from latent space back to the image space (Li et al., 2018). However, these approaches are limited in their performance because each prototype must represent the entire image. ProtoPNet (Chen et al., 2019) introduced the idea of prototypical parts. In this setting, each prototype is a latent patch of the input image, commonly a $1\times 1$ latent patch. The prototypes are each associated with a single class and learned via backpropagation without additional information. The similarity of the prototypes to the image patch is based on the Euclidean distance.
Multiple successors build on the idea of ProtoPNet. ProtoPShare (Rymarczyk et al., 2020) merges the class-exclusive prototypes to class-shared ones. ProtoPool (Rymarczyk et al., 2022) directly learns class-shared patch prototypes and pools them by learning a slot assignment. Deformable ProtoPNet (Donnelly et al., 2021) learns a mixture of prototypical parts with dynamic spatial arrangement. For computing the similarity between the prototypes and image patches, Deformable ProtoPNet uses the cosine similarity. TesNet (Wang et al., 2021) proposes to compute the prototype similarity on the Grassman Manifold.
All these approaches share the use of a linear classification layer. However, ProtoTree (Nauta et al., 2021) builds a Decision Tree on the learned Euclidean prototypes and ProtoKNN (Ukai et al., 2023) proposes to use a $k$-nearest-neighbor classifier with cosine prototype similarities.
Recent work proposes to learn more interesting or robust features for the prototypes from the image encoders. PIPNet (Nauta et al., 2023) use augmentations during training to align the learned prototypes with more meaningfull image content. LucidPPN (Pach et al., 2024) propose a hybrid architecture with texture only grayscale prototypes and low resolution color only prototypes.
@@
Zhou et al. (2022) propose to view prototype learning as a clustering task in the network’s latent space for image segmentation. In this setting, the clustering can be done outside of the network’s backpropagation path with an external clustering algorithm, so-called non-learnable prototypes. Later work such as ProtoGMM (Moradinasab et al., 2024) and MGProto (Wang et al., 2024b) build on this idea and model the prototypes as Gaussian distributions.
ProtoPFormer (Xue et al., 2024) adapts Euclidean prototypical part networks to Transformer based architectures by using a VisionTransformer (ViT, Dosovitskiy, 2020) backbone. To avoid the distraction problem due to the global attention mechanism of ViT architectures, ProtoPFormer uses a hybrid approach with global and local prototypes. The local prototypes are encouraged to focus on different parts of the image by modeling the spatial prototype activations as a 2D Gaussian.
Table 1 presents a systematic overview of the discussed works. The cosine similarity has been shown to perform well in classification tasks (Mettes et al., 2019), but only few prototype learning works have used them so far (Donnelly et al., 2021; Zhou et al., 2022; Ukai et al., 2023). There is also a lack for probabilistic prototypes in a hyperspherical space. We introduce HyperPG, which combines the predictive performance gains from the cosine similarity with the probabilistic nature of Gaussian distributions.
3 Overview: Prototype Formulations
@@
Figure 2: Illustration on how the different prototype formulations compute the similarity between a prototype ${\bm{p}}$ and latent vector ${\bm{z}}$. Euclidean prototypes compute the $L_{2}$ distance between two points in latent space. Hyperspherical prototypes use the cosine similarity of normalized vectors, which corresponds to the angle between two points on a hypersphere. Gaussian prototypes model a Gaussian distribution in Euclidean space and compute the probability density function (PDF). HyperPG prototypes learn a Gaussian distribution of cosine similarities, thereby projecting a Gaussian distribution onto the surface of a hypersphere.
@@
Prototype learning for deep neural networks involves finding structures in latent space representations. This section provides an overview of existing methods, which are also illustrated in Figure 2. Prior work uses point-based prototypes, computing similarity scores relative to a single point in latent space. On the other hand, probabilistic formulations such as Gaussian prototypes or HyperPG prototypes allow the model to adapt to the variance in the training data.
@@ 3.1 Point Based Prototypes
Early approaches based on Autoencoder architectures (Li et al., 2018) use $\rho=\zeta$, meaning the prototype represents the entire image and can therefore be decoded out from latent space back to the input space. Part-based approaches like ProtoPNet and segmentation models like ProtoSeg use $\rho=1$ (Chen et al., 2019; Zhou et al., 2022), meaning each prototype represents some part of the image. Notable exceptions include Deformable ProtoPNet (Donnelly et al., 2021), where each prototype has is a $\rho=3\times 3$ arrangement of smaller patches, and MCPNet (Wang et al., 2024a), which learns concept prototypes across the latent features.
The prediction is computed by comparing each prototype ${\bm{p}}$ to the latent feature map ${\bm{z}}$. For simplicity’s sake lets assume the spatial dimensions $\rho=\zeta=1$. The following equations can be adapted for higher spatial dimensions by summing over the height and width $\sum_{\rho_{w}}\sum_{\rho_{h}}$ for each patch of the latent map.
3.1.1 Euclidean Prototypes
ProtoPNet’s Euclidean prototypes leverage the $L_{2}$ similarity. The $L_{2}$ similarity measure is defined as
@@ 3.1.1 Euclidean Prototypes
and is based on the inverted $L_{2}$ distance between a latent vector ${\bm{z}}$ and a prototype vector ${\bm{p}}$. This similarity measure starts to perform worse with higher numbers of dimensions, as in a large enough space all points are distant to each other and the $L_{2}$ distance looses meaning.
3.1.2 Cosine Prototypes
Prototype formulations based on the cosine similarity create a hyperspherical space. They have been shown to perform well in classification tasks (Mettes et al., 2019) and recent works have started to apply them (e.g., Donnelly et al., 2021; Zhou et al., 2022; Nauta et al., 2021; Ukai et al., 2023). The cosine similarity is defined as
@@ 3.1.2 Cosine Prototypes
which is based on the angle between two normalized vectors of unit length. By normalizing $D$ dimensional vectors to unit length, they are projected onto the surface of a $D$ dimensional hypershere. The cosine similarity is defined on the interval $[-1,1]$ and measures: $1$ for two vectors pointing in the same direction, 00 for orthogonal vectors, and $-1$ for vectors pointing in opposite directions. Without the normalization to unit length, we can define a scaled dot-product prototype inspired by the Attention mechanism (Vaswani et al., 2017). However, this similarity measure is not bounded to any interval, as the vector length is proportionally preserved. For more details see Appendix C. Like the $L_{2}$ similarity, the cosine similarity is a point-based measure comparing two vectors directly.
Both the $L_{2}$ and cosine similarity have been used for classification tasks. The similarity scores are processed by a fully connected layer (e.g., Chen et al., 2019; Donnelly et al., 2021), or a winner-takes-all approach assigns the class of the most similar prototype (e.g., Sacha et al., 2023). The prototypes can be learned by optimizing a task-specific loss, such as cross-entropy, via backpropagation. Alternatively, Zhou et al. (2022) propose “non-learnable” prototypes, whose parameters are obtained via a clustering operation in the latent space instead of backpropagation.
3.2 Probabilistic Prototypes
Viewing prototype learning as clustering problem, point-based prototypes compute the similarity of a latent vector to the center coordinate of each cluster, as represented by the prototype vector. Probabilistic prototypes aim to model the cluster as a probability distribution. Future work could use probabilistic prototypes for extended downstream tasks such as outlier detection or interventions by sampling from generative distributions.
3.2.1 Gaussian Prototypes
Gaussian prototypes model the clusters in latent space as a Gaussian distribution with mean and covariance. By adapting the covariance matrix to the training data a Gaussian prototype with a wide covariance can still have a relatively high response even for larger $L_{2}$ distances from the mean vector. On the other hand, with a very small covariance matrix a Gaussian prototype could show no response unless the mean value is met nearly exactly.
Let the formal definition of a Gaussian prototype be ${\bm{p}}^{G}=(\bm{{\mu}},\bm{\mathrm{\Sigma}})$. The parameters of ${\bm{p}}^{G}_{c,j}$ now track both the mean and covariance of latent vector distribution ${\bm{Z}}_{c,j}$. Each Gaussian prototype ${\bm{p}}^{G}$ thus defines a multivariate Gaussian Distribution ${\mathcal{N}}(\bm{{\mu}},\bm{\mathrm{\Sigma}})$. Gaussian prototypes can be trained using EM for clustering in the latent space (Zhou et al., 2022; Moradinasab et al., 2024; Wang et al., 2024b) or naively by directly optimizing the parameters via Backpropagation. The similarity measure of Gaussian prototypes is defined as the probability density function (PDF) for $D$-dimensional multivariate Gaussians, namely